### v1

# SLA Management in Reconfigurable Multi-Agent RAG: A Systems Approach to Question Answering

Abstract
Retrieval Augmented Generation (RAG) enables Large Language Models (LLMs) to generalize to new information by decoupling reasoning capabilities from static knowledge bases. Traditional RAG enhancements have explored vertical scaling—assigning subtasks to specialized modules—and horizontal scaling—replicating tasks across multiple agents—to improve performance. However, real-world applications impose diverse Service Level Agreements (SLAs) and Quality of Service (QoS) requirements, involving trade-offs among objectives such as reducing cost, ensuring answer quality, and adhering to specific operational constraints.
In this work, we present a systems-oriented approach to multi-agent RAG tailored for real-world Question Answering (QA) applications. By integrating task-specific non-functional requirements — such as answer quality, cost, and latency — into the system, we enable dynamic reconfiguration to meet diverse SLAs. Our method maps these Service Level Objectives (SLOs) to system-level parameters, allowing the generation of optimal results within specified resource constraints.
We conduct a case study in the QA domain, demonstrating how dynamic re-orchestration of a multi-agent RAG system can effectively manage the trade-off between answer quality and cost. By adjusting the system based on query intent and operational conditions, we systematically balance performance and resource utilization. This approach allows the system to meet SLOs for various query types, showcasing its practicality for real-world applications.
Index Terms:
Multi-Agent Systems, Information Retrieval, Service Level Agreement, Quality of Service, Large Language Models, Retrieval Augmented Generation
I Introduction
Question Answering (QA) has been a pivotal area of research since the 1960s, aiming to provide precise answers to natural language questions. Unlike traditional search engines that return a list of hyperlinks to relevant pages, QA systems deliver direct answers to user queries [30].
The advent of generative Large Language Models (LLMs) has led search engines to integrate abstractive QA into their offerings—systems that generate answers which may not exist verbatim in source documents [4]. Additionally, new platforms specializing in abstractive QA, such as Perplexity [20] and ChatGPT Search [19], have emerged.
However, LLM-based abstractive QA exhibits notable limitations, including a lack of generalization to unseen data and a propensity for hallucination, i.e. the generation of false assertions [10]. Retrieval Augmented Generation (RAG) [8] addresses these shortcomings by combining LLMs with external knowledge sources, enabling the LLM to function as a reasoning layer rather than a static repository of facts. RAG retrieves context from external sources such as search engines, traditional databases, vector databases [11], or knowledge graphs [17, 28], and utilizes the LLM for reasoning and response generation.
Enhancements to RAG-based QA have primarily focused on decomposing the task into smaller subtasks and distributing responsibilities among specialized components. While these techniques have improved answer quality, they often introduce additional costs in terms of computation or increased response times.
Real-world QA systems must navigate trade-offs between answer quality and the costs associated with providing those answers. Moreover, they must be dynamically reconfigurable to handle diverse Service Level Agreements (SLAs) and Quality of Service (QoS) requirements arising from various query intents and problem domains. For instance:
  • E-commerce Search: Highly sensitive to response times, where delays can significantly impact user experience and sales conversions [25, 1].
  • Customer Support: Requires not just correct answers but stylistically compliant ones that adhere to the brand-voice guidelines of the domain or company [18].
  • Legal and Healthcare Queries: Demand high answer quality due to the critical nature of the information, with less emphasis on cost or latency constraints [22].
In addition to these domain-specific requirements, real-world RAG systems operate under constantly changing and sometimes adverse conditions. Factors such as network congestion, computational capacity constraints, and security concerns necessitate adaptive strategies. For example, during periods of high network traffic, minimizing external API calls and leveraging local LLM instances can help maintain answer quality while adhering to operational constraints.
By incorporating knowledge about the task domain, QoS requirements, and the operational environment, we can quantify and manage the trade-offs between cost, answer quality, and other non-functional requirements. This enables the design of QA systems that are both efficient and effective across a range of real-world scenarios.
I-A Our Contributions
In this paper, we present a novel system-theoretic framework for SLA management in reconfigurable multi-agent RAG systems, specifically tailored for question-answering applications. Our main contributions are as follows:
  • Dynamically Reconfigurable Horizontal Scaling Framework: We propose a method for horizontally scaling multi-agent RAG systems by replicating agents with diverse configurations. This approach systematically improves performance by enabling the system to dynamically adjust resource allocation based on the specific requirements of each query.
  • Implementation and Experimental Validation: We detail the architecture and implementation of our Abstractive Question Answering Intent Handler. Additionally, we present experimental results that demonstrate the effectiveness of our approach, including comparisons of strategies employed by different agent architectures.
  • Novel Metrics and Dataset Creation: We introduce novel metrics and a dataset creation process to address stylistic adherence requirements, which are critical for many industry users.
By addressing the challenges of SLA management in multi-agent RAG systems, our work contributes to the development of QA systems that are adaptable, efficient, and capable of delivering high-quality answers under diverse operational conditions.
I-B Related Work
Scaling Retrieval-Augmented Generation (RAG) systems has been an active area of research, with most approaches focusing on vertical scaling—decomposing tasks into subtasks assigned to specialized components. Recent studies incorporate Large Language Models (LLMs) at the planning stages to mitigate issues like irrelevant retrieval and focus drift. For instance, PlanRAG uses an LLM to generate retrieval plans, enhancing context relevance [12]. Similarly, Retrieve-Plan-Generate (RPG) separates planning and answering stages to improve long-form question answering [14]. AutoGPT+P introduces planning loops for step-by-step plan generation in dynamic environments [2].
Horizontal scaling through multi-agent systems has also been explored, where multiple agents are synchronously assigned to the same question. LLM-Debate employs multiple LLM agents debating to reach better answers [5]. Chain-of-Thought Self-Consistency (CoT-SC) generates multiple reasoning paths and selects the most consistent answer [27]. Li et al. demonstrate that increasing the number of agents with simple voting improves performance [13].
Temporal scaling techniques involve sequential interactions with LLMs to enhance question-answering performance on complex queries. Shao et al. use iterative querying to refine answers for complex information needs [23]. Query refinement methods, such as RQ-RAG involve rewriting, decomposing, and disambiguating complex queries have also been shown to improve answer quality [3].
While these aforementioned methods enhance answer quality, they often overlook dynamic adaptation to operational constraints critical in real-world applications with strict Service Level Agreements (SLAs).
Gao et al. introduce a modular RAG framework that encompasses horizontal, vertical, and temporal scaling techniques [9]. Zhang et al. propose an agentic information retrieval framework capable of handling complex queries by introducing IR agents with greater levels of autonomy [29]. However, these works lack experimental results on adapting systems to real-world environments.
Arbitration mechanisms have been proposed to improve decision-making in multi-agent systems. Some studies suggest using multiple LLMs to select the best response [26]. These methods improve answer quality but may not address trade-offs such as cost, latency, and security.
Evaluation frameworks like RAGAS assess RAG performance but may omit industry-relevant metrics such as adherence to stylistic constraints [6]. In practical applications, an answer must be not only correct but also stylistically appropriate.
Our work differs by focusing on the dynamic reconfiguration of multi-agent RAG systems to meet diverse SLAs in real-world settings. We integrate task-specific non-functional requirements into system parameters, addressing trade-offs between answer quality, cost, latency, and adherence to stylistic guidelines. This practical framework aims to deploy question-answering systems in industrial contexts with stringent SLA demands.
II Architecture
Our Question Answering System is designed to dynamically reconfigure its components to meet Service Level Agreements (SLAs). The system consists of three primary modules: an Intent Detection Module, a Planning Module, and multiple Intent Handlers. A high-level overview of the architecture is shown in Figure 1.
II-A System Inputs and Outputs
User Query: The system accepts arbitrary input queries from users, such as: ”How do I reset my phone?”, ”How to reset phone”, or ”Burger joints near me”.
These queries vary significantly in intent and complexity, necessitating a robust mechanism for accurate intent recognition and appropriate handling.
Figure 1: A bird’s eye view of the architecture of the AbstractiveQA System
Task SLA/QoS Parameters: Specific requirements such as:
  • Required precision and recall for the task
  • Maximum acceptable hallucination rate and incongruent response rate (to ensure adherence to stylistic constraints)
  • Budget and Latency constraints for different intent types
Operational Environment Current conditions affecting the system, such as:
  • Availability of network resources
  • Status of external APIs and services
  • Available models and data sources
The QA system’s outputs are Intent-Dependent Answers which consist of diverse responses tailored to the user’s intent and complying with specified SLAs.
II-B Intent Detection Module
The Intent Detection Module processes the user query to classify it into predefined intent categories, which may include:
  • Directly Answerable Question: Queries seeking specific information that can be directly addressed.
  • Request for Summarization: Queries asking to condense or summarize information from a source.
  • Non-Question Statements: Inputs that are not interrogative but may require action.
  • Request for a List: Queries seeking a collection of items (e.g., nearby restaurants serving hamburgers).
  • Sales Inquiry: Queries related to purchasing or product information.
II-B1 Implementation Options
Implementation options for this module include various techniques, each with trade-offs in terms of computational cost, extensibility, and accuracy:
  • Heuristic Methods: Rule-based systems using predefined patterns and keywords for quick intent classification. While low in computational overhead, they may lack flexibility and scalability.
  • Machine Learning Classifiers: Lightweight models trained on labeled datasets to predict intents, offering a balance between performance and resource utilization.
  • Generative Large Language Models (LLMs): Zero-shot or few-shot classifiers leveraging pre-trained LLMs to understand and classify intents without extensive domain-specific training data. Although highly extensible and accurate, they incur higher computational costs.
II-C Planning Module
The Planning Module dynamically configures the resources of the Intent Handlers to meet the system’s SLA requirements for each specific intent in the current environment. Its core functions include:
  • Resource Allocation: Determining the optimal number of Retrieval Augmented Generation (RAG) agents to deploy, balancing response quality against computational cost.
  • Backend Data Source Selection: Choosing appropriate data repositories (e.g., cached datasets, real-time databases, knowledge graphs) based on required information freshness and relevance.
  • Arbitration Mechanism Configuration: Selecting and parameterizing arbitration algorithms to aggregate and evaluate responses from multiple agents effectively.
For example, to service a query requiring high answer quality, the Planning Module may:
  • Increase Agent Replication: Deploy more RAG agents to generate diverse candidate answers.
  • Employ Appropriate Data Sources: Access multiple data repositories to enrich the context and determine which agents should have access to which sources e.g. an agent using a third-party LLM would not get access to sensitive documents.
  • Select what threshold to use in a voting-based arbitration mechanism e.g. only return an answer when 70% of the agents think it should be answered.
By adjusting these configurations in real-time, the Planning Module ensures compliance with SLA parameters, optimizing the system’s performance in terms of answer quality and resource utilization.
Figure 2: Possible Instantiation of Abstractive Question Answer Intent Handler
II-D Intent Handlers
The Intent Handler is a reconfigurable module that executes the necessary processes to address the user’s query based on the classified intent. By dynamically adjusting its internal workflows and resource utilization in response to directives from the Planning Module, the Intent Handler can:
  • Manage Computational Costs: Allocate resources judiciously to stay within budgetary constraints without significantly compromising answer quality.
  • Enhance Answer Quality: Allocate additional resources, such as more RAG agents or advanced arbitration mechanisms, for queries where accuracy is paramount.
  • Ensure Compliance with Non-Functional Requirements: Adjust processing to adhere to stylistic guidelines, security policies, and other non-functional requirements specified in the SLAs.
To achieve these objectives, the Intent Handler orchestrates a network of logical entities and services, which may include:
  • Backend Data Access Interfaces: APIs and connectors to interact with various data sources.
  • RAG Agents: Multiple instances of Retrieval Augmented Generation agents, each potentially configured with different retrieval strategies, preprocessing pipelines, and LLMs to generate diverse candidate responses.
  • Arbitration Algorithms: Mechanisms to evaluate and select the optimal response from the candidate answers. This may involve:
    • Voting Systems: Methods like majority voting or weighted voting based on agent confidence scores.
    • Ranking Models: Machine learning models trained to rank candidate responses based on relevance, accuracy, and other quality metrics.
  • Ensemble Techniques: Combining outputs from multiple agents to produce a more robust final answer.
  • Response Synthesis: Constructing the final response, ensuring coherence and adherence to stylistic guidelines.
A diagram showing a potential instantiation of a multi-agent Abstractive QA Intent Handler is presented in Figure 2.
III Benchmark Dataset Creation and Metrics
III-A Dataset Creation
We sampled approximately 5,000 actual user queries spanning over 200 different businesses across various industries (e.g. retail, finance, etc.) Each query was classified into one of two intent categories: Directly Answerable Questions or All Other Intents.
In this context, a query is considered directly answerable if it can be addressed succinctly (e.g., with a yes/no or factual response), such as inquiries about a product, service, or specific company information. Non-directly answerable queries include broad topics, technical issues, or those better answered with a list or document link.
Each query was annotated by two independent annotators, followed by a conflict resolution process when necessary. The resulting dataset served two purposes:
  • Evaluation of the Intent Classifier (which is outside the scope of this paper).
  • Creation of an annotated dataset for queries and answers.
We found that directly answerable queries constituted approximately 12% of the original dataset, totaling around 600 queries. Each of these queries was input into five separately configured RAG-QA Agents, resulting in approximately 3,000 (query, context, answer) triples. Each triple was annotated by two independent annotators, with a conflict resolution step to ensure consistency.
Each result was assigned with a total of three tags dependent on the availability of an answer in context, the provision of an answer, and the quality of the answer: Context Availability:
  • Answer Exists in Context: The agent’s retrieval process returned at least one result from which a correct direct answer could be sourced.
  • Answer Does Not Exist in Context: The agent’s retrieval process did not return any results from which a correct direct answer could be sourced.
Answer Provision:
  • Answer Provided: An answer was returned.
  • No Answer Provided: The returned answer was null.
Answer Quality:
  • Correct Answer: The returned answer was correct and aligned with the task’s non-functional requirements (e.g., adhered to the specified style in the prompt).
  • Hallucination: The returned answer was not grounded in the provided context.
  • Incongruent Response: The returned answer did not align with the task’s non-functional requirements.
III-B Metrics
III-B1 Answer Quality Metrics
We evaluated the performance of the agents using the following metrics:
  • Precision: The number of correct answers divided by the total number of answers provided.
  • Recall: The number of correct answers divided by the number of times the answer exists in the global context (the union of all contexts from all agents). [^1] The use of the global context to compute precision and recall introduces limitations to our dataset, as the exact upper bound on the number of answers in any context is unknown and could not be determined within our time constraints. However, we found that the difference in the number of answers sourced from the global context between the four-agent and five-agent systems was minimal (one to two instances, depending on configuration). This suggests that the overall impact is likely negligible with the retrieval mechanisms used.
  • F1 Score: The harmonic mean of precision and recall.
    $F_{1}=2\cdot\frac{\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$
  • Hallucination Rate: The number of hallucinations divided by the total number of answers provided, where a hallucination is any answer not grounded in the context.  [^2] The Hallucination Rate is analogous to the False Positive Rate, with the key difference being that false positives in this context include answers not grounded in the provided context (extrinsic hallucinations), contradicted by the provided context (intrinsic hallucinations), or incongruent responses. Further details on categories of hallucination can be found in [10].
  • Incongruent Response Rate: The number of incongruent responses divided by the total number of responses.
Precision, Recall, and F1: In traditional information retrieval, precision is defined as the proportion of retrieved documents that are relevant, and recall is the proportion of relevant documents that are retrieved [16]. In our question answering domain, we adapted these definitions to suit our specific context.
Incongruent Responses: For many users of this system, particularly brands, it is insufficient to simply return a correct answer; the answer must also adhere to the non-functional and stylistic requirements specified in the prompt. This includes guidelines such as not breaking the fourth wall or referencing the context directly. Ensuring compliance with these requirements is essential for maintaining brand voice and user experience.
III-B2 System Cost
The expected cost of our QA system, as used in our experiments, is calculated as the sum of the fixed system overhead and the costs associated with QA agents and arbitration, aggregated across all intents. The formula is:
$C_{\text{sys}}=C_{\text{overhead}}+\sum_{\text{intent}=1}^{I}\big{(}C_{\text{agent}}\cdot N_{\text{intent}}+C_{\text{arbitration}}(N_{\text{intent}})\big{)}$
Where:
  • $C_{\text{sys}}$: Total expected system cost.
  • $C_{\text{overhead}}$: Fixed system overhead cost.
  • $C_{\text{agent}}$: Cost per QA agent.
  • $N_{\text{intent}}$: Number of agents for a specific intent.
  • $C_{\text{arbitration}}(N_{\text{intent}})$: Arbitration cost for $N_{\text{intent}}$ agents.
  • $I$: Total number of intents.
This equation models the specific cost structure of our QA system, where both the agent and arbitration costs are functions of the number of agents configured for each intent.
III-B3 System Latency
The system latency for our implementation is defined as the sum of the fixed overhead latency, the maximum latency of the QA agents configured for a given intent (since agent operations are asynchronous), and the arbitration latency for those agents. Mathematically, this is expressed as:
$L_{\text{sys}}=L_{\text{overhead}}+\max_{i\in\text{agents}}(L_{\text{QA},i})+L_{\text{arbitration}}(N_{\text{intent}})$
Where:
  • $L_{\text{sys}}$: Total system latency.
  • $L_{\text{overhead}}$: Fixed overhead latency.
  • $L_{\text{QA},i}$: Latency of the $i$-th QA agent.
  • $\max_{i\in\text{agents}}(L_{\text{QA},i})$: Maximum latency among all QA agents included in the intent service.
  • $L_{\text{arbitration}}(N_{\text{intent}})$: Arbitration latency for $N_{\text{intent}}$ agents.
This formula reflects the asynchronous nature of the QA agent operations and the specific latency structure of our experimental implementation.
IV Agent Implementation
In this section, we describe the experiments conducted to evaluate our reconfigurable multi-agent RAG system for question answering. We detail the agent architecture, single-agent experiments, and the arbitration algorithm, including its implementation and ablation studies.
Figure 3: The RAG QA Agent Architecture we use in our experiments
IV-A Agent Architecture
The architecture of an individual QA agent used in our experiments, shown in Figure 3, is designed to be flexible and configurable based on the output of the Planning Module. The specific parameterization of each agent depends on factors such as which backend data sources are provisioned and caching strategies. While the implementation details of the Planning Module are outside the scope of this paper, we provide an overview of the QA agent’s architecture and workflow.
  • Query Input: The user query is input into the Retrieval Module. Only queries classified as directly answerable are processed by the QA agent.
  • Retrieval Module: Accesses backend data sources using the input query. Utilizes a federated search approach, querying multiple backend data sources asynchronously. Supports vertical search [^3] Vertical search refers to specialized searches within predefined categories or domains. For instance, a healthcare-related query for cardiology may return results segmented into verticals such as cardiologists and hospitals specializing in cardiology. by returning data categorized by entity type, configured based on domain requirements. Verticals are ranked in order of relevance.
  • Post-Retrieval Preprocessing: A preprocessing module prepares the context documents for input into the Reasoning Module. Prunes and organizes retrieved documents to comply with the LLM’s context window size limitations.
  • Reasoning Module: Powered by GPT-4 [7]. Generates answers based on the user query and preprocessed context. Details of the prompt used are discussed in Section IV-C.
Figure 4: Architecture of the Vector Search Process we use in our experiments
IV-B Retrieval Module Details
The Retrieval Module, shown in Figure 4, pulls data from various backend sources, including vector databases, lexical search engines, and graph data stores, depending on the domain. For vector search, we employ the following components:
  • Embedding Model: Documents are embedded using a transformer model based on MPNet [24]. Embeddings capture semantic information for effective similarity comparisons.
  • Vector Database: Indexed using Hierarchical Navigable Small World (HNSW) graphs [15] for efficient approximate nearest neighbor search. At query time, the same embedding model is used to encode the user query. The resulting query vector $\mathbf{q}$ is used to retrieve documents whose vectors $\mathbf{d}$ are similar, based on cosine similarity:
    $similarity(\mathbf{q},\mathbf{d})=\frac{\mathbf{q}\cdot\mathbf{d}}{\|\mathbf{q}\|\|\mathbf{d}\|}$
    Documents with higher cosine similarity scores are considered more relevant to the query.
Prompt Example prompt = f""" Context: {context_entities}. Use only the provided context to answer the question: "{query}" to the best of your ability and in a few sentences or less. If there is more than one answer, summarize the options. Provide the "uid" values of the object you used to inform the answer (do NOT use the "id" value) Return the answer in a json dictionary format [Example: {"answer":"This is the answer", "uid_list": [12345, 98342]}] Do not refer to the context in your answer. """
Figure 5: An example of the constructed prompt used by the QA Agent’s Reasoning Module.
IV-C Prompt Engineering
After re-ranking and/or thresholding, we construct the prompt for the LLM. The preprocessed list of context documents and the original user query are interpolated into the prompt at appropriate places. An example prompt is shown in Figure 5:
V Experiments
We conducted experiments on individual QA agents by varying the post-retrieval preprocessing logic used to prepare the input prompts for the GPT-4-powered Reasoning Module. Given the LLM’s context window limitation of 8,000 tokens, we needed to prune the retrieved context appropriately. We developed several pruning methods aiming to balance content relevance with the token limit. For each agent, we measured Answer Quality using the metrics described in Section III-B1.
For the multi-agent experiments, we implemented an arbitration algorithm based on majority voting and relevance ranking. We varied the size of the agent ensemble dedicated to servicing the QA queries by adjusting the number of QA agents—each employing diverse post-retrieval preprocessing strategies—included in the ensemble. The final answer was selected based on the arbitration algorithm. Similar to the single-agent experiments, we measured Answer Quality and evaluated the system cost relative to the number of agents in the ensemble. [^4] A notable limitation of these experiments was the inability to accurately measure system latency due to non-stationary external conditions, such as API rate limiting. Future work will involve simulating realistic conditions to accurately measure latency.
V-A Preprocessing Methods
We implemented two main types of preprocessing logic:
  • Thresholding:
    • Thresholding (Control): A naive truncation strategy where the first 8,000 tokens from the search results are taken.
    • Vertical Thresholding: Selects top tokens from the top two verticals (until the prompt reaches 8,000 tokens).
    • Aggressive Thresholding: Takes the first 6,000 tokens from the search response, allowing more space for the reasoning process.
  • Re-ranking: Utilized a cross-encoder model based on ’ms-marco-MiniLM-L-12-v2’ from the Sentence Transformers library [21]. The cross-encoder accepts the query and document as input and returns a relevance score between 0 and 1. Documents are re-ranked based on relevance scores before thresholding.  [^5] Note: The raw entity JSON is not used directly as input. Search configurations define an LLM-friendly view of the entity, specifying which fields should be included. The design of this view generation model is outside the scope of this work.
V-B Arbitration
In our multi-agent RAG system, we employ an arbitration algorithm to select the optimal response from a set of candidate answers generated by various QA agents. The arbitration process is governed by a decision threshold $T\in(0,1)$ and an arbitration function $A$ that maps a set of results to a single final answer. The algorithm is defined in Algorithm 1
Input: Set of candidate responses $C=\{c_{1},c_{2},\dots,c_{n}\}$; threshold $T$
Output: Final answer or a negative result
Step 1: Candidate Response Collection
Let $C$ be the set of candidate responses from $n$ QA agents.
Step 2: Affirmative Response Identification
Define $C_{\text{affirmative}}\subseteq C$ as the subset of responses that include an attempted answer (i.e., non-null and substantive responses).
Step 3: Threshold Evaluation
Compute the minimum number of affirmative responses required:
$k=\left\lfloor T\times|C|\right\rfloor$
if $|C_{\text{affirmative}}|\geq k$ then
       Proceed to arbitration.
else
       Return a negative result indicating that no answer will be provided.
       return Negative Result
Step 4: Arbitration Function Application
Apply the arbitration function $A$ to $C_{\text{affirmative}}$:
Final $\text{Final Answer}=A(C_{\text{affirmative}})$
return Final Answer
Algorithm 1 Candidate Response Collection and Arbitration
For our experiments, the arbitration function $A$ was implemented using a cross-encoder model, consistent with the post-retrieval preprocessing methods described earlier. The cross-encoder jointly encodes each candidate answer with the original query to compute a relevance score. It effectively re-ranks the affirmative responses based on their contextual relevance and quality, selecting the top-ranked answer as the final output.
V-C Algorithm Ablations
To evaluate the effectiveness of our arbitration algorithm and understand the impact of its components, we conducted a series of ablation experiments:
  • Random Candidate Selection:
    • Method: Select a candidate answer at random from the set of affirmative responses, bypassing the arbitration function.
    • Purpose: Establish a baseline to assess the added value of the arbitration function in selecting higher-quality answers compared to random selection.
  • Voting Procedure with Random Positive Selection:
    • Method: Apply the threshold evaluation to determine if sufficient affirmative responses exist. If so, randomly select one of the affirmative answers without applying $A$.
    • Purpose: Isolate the effect of the threshold-based decision while assessing the necessity of the arbitration function for answer selection.
  • Cross-Encoder Arbitration on All Candidate Responses:
    • Method: Apply the cross-encoder-based arbitration function $A$ to all candidate responses, including both affirmative and negative ones (e.g., null or placeholder answers).
    • Purpose: Examine whether including all candidate responses in the arbitration process affects the quality of the final answer and whether the arbitration function can effectively discriminate among all types of responses.
These ablation experiments allowed us to analyze the contributions of different components of the arbitration algorithm:
  • Effectiveness of the Arbitration Function $A$:
    • By comparing the random selection methods with the cross-encoder arbitration, we assessed how much the arbitration function improves answer quality by selecting the most relevant response.
  • Impact of the Decision Threshold $T$:
    • Evaluating different values of $T$ helped determine the optimal balance between answer availability and quality.
    • A higher threshold requires more consensus among agents before providing an answer, potentially reducing incorrect or low-quality responses.
  • Inclusion of Negative Responses:
    • Testing the arbitration function on all candidate responses explored whether the model could effectively handle and filter out non-informative or irrelevant answers.
VI Results
In this section, we present the results of our experiments, focusing on both single-agent ablations and multi-agent configurations. Our goal is to evaluate how different preprocessing strategies and arbitration mechanisms affect the performance of the QA system, particularly in terms of precision, recall, hallucination rate, and incongruent response rate.
VI-A Single-Agent Experiments
The results of the single-agent experiments are summarized in the upper section of Table I. These results encompass the three preprocessing strategies discussed in the Single Agent Experiments section
Key Findings
  • Aggressive Thresholding emerged as the best-performing strategy, achieving the highest precision and recall, as well as the lowest incongruent response rate.
  • Vertical Thresholding performed comparably to aggressive thresholding, indicating that focusing on top-ranked verticals effectively preserves relevant context.
  • Agents utilizing Re-ranking with Cross-Encoder exhibited lower precision and recall. The cross-encoder tended to eliminate important context documents, leading to reduced precision despite a lower hallucination rate.
TABLE I: Experimental Results for Single and Multi Agent Experiments
Experiment Version Recall Precision F1 Hallucination Rate Incongruent Response Rate
Thresholding (control) 0.640 0.656 0.648 0.239 0.015
Rerank with thresholding 0.529 0.652 0.584 0.187 0.027
Rerank with vertical thresholding 0.571 0.619 0.594 0.249 0.017
Aggressive Thresholding 0.656 0.670 0.663 0.235 0.008
Vertical Thresholding 0.656 0.654 0.655 0.249 0.013
vote_with_thresh, N=3 0.655 0.672 0.663 0.230 0.011
vote_with_thresh, N=5 0.684 0.691 0.688 0.220 0.012
vote_most_relevant, N=3 0.666 0.674 0.670 0.229 0.014
vote_most_relevant, N=5 0.684 0.681 0.683 0.229 0.013
vote_most_relevant_with_thresh, N=3 0.652 0.672 0.661 0.228 0.013
vote_most_relevant_with_thresh, N=5 0.684 0.691 0.688 0.220 0.012
VI-B Multi-Agent Experiments
The multi-agent experiments were designed to evaluate the impact of ensemble configurations on system performance. We varied the number of agents in the ensemble and observed the effects of different arbitration mechanisms, all of which used a threshold $T=0.5$, requiring at least half of the agents to provide an affirmative answer for a response to be returned. The results can be found in the bottom section of Table I. Detailed charts are shown in Figures 6 - 10.
Key Findings
  • Increasing the number of agents consistently improved performance, enhancing precision and recall while reducing the hallucination and incongruent response rates. This highlights the benefit of aggregating responses across agents.
  • Balancing performance improvements with computational costs is crucial. Larger ensembles offer better performance metrics but come at the expense of increased resource usage and latency.
  • Simple voting mechanisms often performed on par with relevance-based arbitration. While the latter showed predictable improvements with more agents, the additional gains were minimal compared to the increased computational overhead.
Figure 6: Recall vs. Ensemble Size
Figure 7: Precision vs. Ensemble Size
Figure 8: F1 vs. Ensemble Size
Figure 9: Hallucination Rate vs. Ensemble Size
Figure 10: Incongruent Response Rate vs. Ensemble Size
VII Discussion
Our experiments demonstrate the effectiveness of dynamically reconfigurable multi-agent RAG systems in managing trade-offs between answer quality and resource utilization. Insights Effectiveness of Aggressive Thresholding: By aggressively limiting the context to the most relevant information, we reduce noise and enhance the LLM’s ability to generate accurate answers. Limitations of Re-ranking Alone: While re-ranking aims to prioritize relevant documents, it can inadvertently discard essential context, adversely affecting performance. Scalability of Multi-Agent Systems: Adding more agents generally improves performance but increases cost and latency. Systems must be designed to adjust dynamically based on SLA requirements.
VII-A Future Work
  • Exploration of Additional Arbitration Mechanisms: Investigate other methods, such as weighted voting or machine learning-based arbitration, to enhance decision-making.
  • Incorporation of QoS Parameters: Extend the framework to include latency and computational constraints explicitly, enabling more granular SLA management.
  • Complexity of Planning Mechanisms: Develop more sophisticated Planning Modules capable of temporal replication (e.g., handling requests for refinement) and adjusting strategies in real-time.
  • Simulation of Real-World Conditions: Incorporate factors like network congestion and varying computational capacities to assess the system’s robustness under diverse operational environments.
  • Threshold Parameter Optimization: Study the impact of different threshold values in arbitration mechanisms to find optimal settings for various use cases.
Our case study illustrates that a systems-oriented approach to SLA management in reconfigurable multi-agent RAG systems is both feasible and beneficial. By dynamically adjusting system parameters based on non-functional requirements, we can effectively balance answer quality with resource constraints. This work serves as a foundational step toward deploying RAG-based QA systems in real-world applications with diverse and stringent SLA demands.